1. Convert the decimal number +164 and -164 to 9-bit binary numbers according to Sign magnitude, One’s complement, and Two’s complement

2. Convert the binary number 111011010 to base 10 decimal form (our regular number system) treating it as each of the following representations: Sign magnitude, One’s complement, and Two’s complement
Answer #1

2) Given 1 1101 1010

In Sign magnitude the

The sign "1" means negative

11011010 = (1 × 2⁸) + (1 × 2⁷) + (0 × 2⁶) + (1 × 2⁵) + (1 × 2⁴) + (0 × 2³) + (1 × 2²) + (0 × 2¹) + (-16 × 2⁰) = 218

The magnitude is 218

In One’s complement

111011010 = (1 × 2⁸) + (1 × 2⁷) + (1 × 2⁶) + (0 × 2⁵) + (1 × 2⁴) + (1 × 2³) + (0 × 2²) + (1 × 2¹) + (0 × 2⁰) = 474

In Two’s complement

111011010 = -38